OpenStudy (jazzyfa30):

Solve the equation by completing the square: x^2 + 2x = 120

OpenStudy (anonymous):

@jazzyfa30 DO U KNW HW TO SOLVE A QUADRATIC EQUATION USING THE METHOD OF COMPLETING SQUARE???

OpenStudy (jazzyfa30):

no

OpenStudy (anonymous):

DO U KNW THE ALZEBRAIC IDENTITY (A+B)^2=A^2+2AB+B^2

OpenStudy (jazzyfa30):

no

OpenStudy (anonymous):

so let us first discuss the alzebraic identity (a+B)^2=a^2+2ab+b^2

OpenStudy (anonymous):

let us the find the square of 125=125^2=??? nw let us apply the identity here we can split 125 as 100+25 125^2=(100+25)^2

OpenStudy (anonymous):

nw in the above example a=100 b=25 nw a^2+2ab+b^2=100^+2.100.25+25^2 =10000+5000+625=15625 threfore 125^2=15625

OpenStudy (jazzyfa30):

what do u have a skype

OpenStudy (anonymous):

multiplying 125 and 125 will take a lot of time but applying the identity in the example will of course save some time

OpenStudy (anonymous):

have u understood the identity @jazzyfa30??? if yes then let's come to ur question

OpenStudy (jazzyfa30):

no thats y i asked u whether u had a skype so u could explain it verbally cus im not a visible learner

OpenStudy (anonymous):

x=-12,10

OpenStudy (jazzyfa30):

@pascuzzoa please dont give me the answer

OpenStudy (anonymous):

oops, well its to late now

OpenStudy (anonymous):